OpenStudy (anonymous):

Find the value of √(4^2 + 3^2 - 10^2)

OpenStudy (anonymous):

\[\sqrt{16+9-100}\]is a start

OpenStudy (anonymous):

i get \(\sqrt{-75}\) which is not a real number, but you can write it as \[5\sqrt{3}i\] if you like

OpenStudy (anonymous):

Thanks